# Computing Distances on Graph Associahedra is Fixed-parameter Tractable

### v1

@@ 4.4 Restricting the rotations to marked vertices
Note that, if in the statement of 4.14 we replaced “at most $k$ vertices” with “at most $2k$ vertices”, then its proof would be trivial, as any of the at most $k$ rotations of $\sigma$ involves two vertices, so at most $2k$ distinct vertices overall. In that case, for the proof of 4.16 to go through, we would have to replace, in Equation 4 in the definition of type, “$k+1$” with “$12k+$ 1” when taking the minimum. In the sequel we will often use a weaker version of 4.14, namely that for every vertex $v\in V(T)$, at most $k$ vertices in ${\sf children}(T,v)$ are used by an $\ell$-rotation sequence from $T$ to $T^{\prime}$.
@@
To this end, let $v\in V(T)\setminus M$ be a furthest (with respect to the distance to ${\sf root}(T)$) non-marked vertex of $T$ that is used by $\sigma$. By 4.3, we can assume that $v\in B_{\sf cb}$. Let $Z$ be the connected component of $T[B_{\sf cb}]$ such that $v\in Z$. For technical reasons, it will be helpful to assume that $v$ is not a leaf of $Z$. (This can be achieved, for instance, by observing that the analysis of the size of the components $Z$ in 4.3 is not tight. Alternatively, we can just “artificially” increase their diameter by one –i.e., replacing $2k$ with $12k+$ 1 in the definition of $B_{\sf cb}$– so that we can safely assume that the leaves of $Z$ are never used by a rotation sequence.) Note that the choice of $v$ as a lowest (i.e., furthest) non-marked vertex of $T$ used by $\sigma$ implies that for every vertex $u\in{\sf children}(T,v)$, the whole subtree $T(u)$ remains intact throughout $\sigma$, meaning that it appears as a rooted subtree in all the intermediate elimination trees generated by the $\ell$-rotation sequence $\sigma$. We distinguish two cases, the second one being considerably more involved, but that will benefit from the intuition developed in the first one.
@@
We need to prove that if $\sigma=(e_{1},\ldots,e_{\ell})$ and $e_{i}=vw$ for some $i\in[\ell]$ and some $w\in V(T)$, then $v^{\prime}w\in E(T^{\prime}_{i-1})$. Let us prove it by induction. For $i=1$, assume that $e_{1}=vw$ for some $w\in V(T)$. The choice of $v$ as a lowest non-marked vertex of $T$ used by $\sigma$ implies that $w={\sf parent}(T,v)$, and since $v^{\prime}$ is a $T$-sibling of $v$, it follows that $v^{\prime}w\in E(T^{\prime}_{0})=E(T)$. Assume now inductively that the edges to be rotated exist up to $i-1$, and suppose that $e_{i+1}=vw$ for some $w\in V(T)$. Since $\tau(T,Z,v)=\tau(T,Z,v^{\prime})$, it follows in particular that ${\sf trace}(T,Z,v)={\sf trace}(T,Z,v^{\prime})$ (see the green dotted edges in Figure 8), and as $v$ and $v^{\prime}$ are $T$-siblings, it follows that the rooted subtrees $T(v)$ and $T(v^{\prime})$ have exactly the same neighbors in the set $V(Z)\setminus(V(T(v))\cup V(T(v^{\prime})))$, that is,
@@
Equation 8 implies that, up to $i-1$, vertex $v^{\prime}$ has been following exactly the same moves in $\sigma^{\prime}$ as the ones followed by vertex $v$ in $\sigma$. Thus, the fact that $vw\in E(T_{i})$ implies that $v^{\prime}w\in E(T^{\prime}_{i})$, and the claim follows.
@@
Since ${\sf root}(Z)\in M$ and $v\notin M$, the definition of the marking algorithm implies that there exists a marked $T$-sibling $v^{\prime}$ of $v^{\star}$ with $\tau(T,Z,v^{\star})=\tau(T,Z,v^{\prime})$. Moreover, the fact that $M$ is defined by recursively marking up to $k+1$ vertices of each type implies, together with 4.14 and the fact that the desired vertex $v^{\star}$ is non-marked, that there is some $T$-sibling $v^{\prime}$ of $v^{\star}$ with $\tau(T,Z,v^{\star})=\tau(T,Z,v^{\prime})$ such that no vertex in $T(v^{\prime})$ is used by $\sigma$. Finally, we can clearly choose $v^{\star}$ in a unique way as being the vertex closest to $v$ satisfying the above properties.
Figure 9: Illustration of Case 2 in the proof of 4.16. All vertices in $T(v^{\star})$ are non-marked, and (at least) vertex $v$ is used by $\sigma$. No vertex in $T(v^{\prime})$ is used by $\sigma$, and (at least) vertex $v^{\prime}$ is marked (in red). The sets $U_{i,\tau}$ of vertices used by $\sigma$ in $T(v^{\star})\cap Z$ are depicted with squares, as well as their images in $T(v^{\prime})\cap Z$ via the representative function $\rho$.
@@
Assuming that the sets $U^{\prime}_{i,\tau}$ exist, it is clear that $\rho$ is injective. Hence, we shall prove that for every $i=0,\ldots,{\sf dist}_{T}(v^{\star},v)$ and every type $\tau$ occurring in a vertex in $U_{i}$, there exists a set $U^{\prime}_{i,\tau}\subseteq V(T(v^{\prime}))\cap Z$ of marked vertices of type $\tau$ with $|U^{\prime}_{i,\tau}|=|U_{i,\tau}|$. We proceed by induction on $i$. For $i=0$, $U_{0}=\{v^{\star}\}$ and the only type occurring in $U_{0}$ is $\tau(T,Z,v^{\star})=:\tau$. Thus, we can just take $U^{\prime}_{0,\tau}=\{v^{\prime}\}$, where $v^{\prime}$ is the vertex given by 6, so that $\rho(v^{\star})=v^{\prime}$. We now prove the statement for any $i\geq 1$, using the recursive definition of type (cf. 4.9). We may assume that $U_{i}\neq\emptyset$, as otherwise there is nothing to prove. The facts that $\tau(T,Z,v^{\star})=\tau(T,Z,v^{\prime})$ and that $v^{\prime}$ is marked imply that, for every $i\in[{\sf dist}_{T}(v^{\star},v)]$ and every type $\tau$, the following holds:
@@
  • Suppose first that none of $u,w$ belongs to $T(v^{\star})$. It is not difficult to verify that if $uw\in E(T_{i-1})$, then $uw\in E(T^{\prime}_{i-1})$ as well, where $T^{\prime}_{i-1}$ is the tree obtained from $T$ by applying the first $i-1$ rotations of $\sigma^{\prime}$. Indeed, the fact that $\tau(T,Z,v^{\star})=\tau(T,Z,v^{\prime})$ implies that the existence of such an edge with both endpoints outside $T(v^{\star})$ is preserved when replacing the vertices in $U$ with their representatives.
@@
    Assume now inductively that the edges to be rotated exist up to $i-1$, and suppose that $e_{i+1}=uw$ for some $u\in V(T(v^{\star}))$ and $w\notin V(T(v^{\star}))$. The fact that $\tau(T,Z,u)=\tau(T,Z,\rho(u))$ implies that ${\sf trace}(T,Z,u)={\sf trace}(T,Z,\rho(u))$, which implies in particular that $T(u)$ and $T(\rho(u))$ have exactly the same neighbors in the set $V(Z)\setminus(V(T(v^{\star}))\cup V(T(v^{\prime})))$, that is,
@@
    The fact that $\tau(T,Z,u)=\tau(T,Z,\rho(u))$ and Equation 9 imply that, up to $i-1$, vertex $\rho(u)$ has been following the same moves within $T(v^{\prime})$ and $V(Z)\setminus(V(T(v^{\star}))\cup V(T(v^{\prime})))$ in $\sigma^{\prime}$ as the ones followed by vertex $u$ within $T(v^{\star})$ and $V(Z)\setminus(V(T(v^{\star}))\cup V(T(v^{\prime})))$ in $\sigma$. Thus, the fact that $uw\in E(T_{i})$ implies that $\rho(u)w\in E(T^{\prime}_{i})$, and the claim follows.
@@
By 9, $\sigma^{\prime}$ is an $\ell$-rotation sequence from $T$ to $T^{\prime}$, and it uses strictly more marked vertices than $\sigma$, because no vertex of $T(v^{\star})$ is marked (by the conditions in 6), and within $T(v^{\prime})$ there is at least one marked vertex, namely $v^{\prime}$. This concludes Case 2.

### v2

@@ 4.4 Restricting the rotations to marked vertices
Note that, if in the statement of 4.14 we replaced “at most $k$ vertices” with “at most $2k$ vertices”, then its proof would be trivial, as any of the at most $k$ rotations of $\sigma$ involves two vertices, so at most $2k$ distinct vertices overall. In that case, for the proof of 4.16 to go through, we would have to replace, in Equation 4 in the definition of type, “k+1k+1” with “2k+12k+1” when taking the minimum. In the sequel we will often use a weaker version of 4.14, namely that for every vertex $v\in V(T)$, at most $k$ vertices in ${\sf children}(T,v)$ are used by an $\ell$-rotation sequence from $T$ to $T^{\prime}$.
@@
To this end, let $v\in V(T)\setminus M$ be a furthest (with respect to the distance to ${\sf root}(T)$) non-marked vertex of $T$ that is used by $\sigma$. By 4.3, we can assume that $v\in B_{\sf cb}$. Let $Z$ be the connected component of $T[B_{\sf cb}]$ such that $v\in Z$. For technical reasons, it will be helpful to assume that $v$ is not a leaf of $Z$. (This can be achieved, for instance, by observing that the analysis of the size of the components $Z$ in 4.3 is not tight. Alternatively, we can just “artificially” increase their diameter by one –i.e., replacing $2k$ with $2k+1$ in the definition of $B_{\sf cb}$– so that we can safely assume that the leaves of $Z$ are never used by a rotation sequence.) Note that the choice of $v$ as a lowest (i.e., furthest) non-marked vertex of $T$ used by $\sigma$ implies that for every vertex $u\in{\sf children}(T,v)$, the whole subtree $T(u)$ remains intact throughout $\sigma$, meaning that it appears as a rooted subtree in all the intermediate elimination trees generated by the $\ell$-rotation sequence $\sigma$. We distinguish two cases, the second one being considerably more involved, but that will benefit from the intuition developed in the first one.
@@
We need to prove that if $\sigma=(e_{1},\ldots,e_{\ell})$ and $e_{i}=vw$ for some $i\in[\ell]$ and some $w\in V(T)$, then $v^{\prime}w\in E(T^{\prime}_{i-1})$. Let us prove it by induction. For $i=1$, assume that $e_{1}=vw$ for some $w\in V(T)$. The choice of $v$ as a lowest non-marked vertex of $T$ used by $\sigma$ implies that $w={\sf parent}(T,v)$, and since $v^{\prime}$ is a $T$-sibling of $v$, it follows that $v^{\prime}w\in E(T^{\prime}_{0})=E(T)$. Assume now inductively that the edges to be rotated exist up to i1i-1, and suppose that $e_{i+1}=vw$ for some $w\in V(T)$. Since $\tau(T,Z,v)=\tau(T,Z,v^{\prime})$, it follows in particular that ${\sf trace}(T,Z,v)={\sf trace}(T,Z,v^{\prime})$ (see the green dotted edges in Figure 8), and as $v$ and $v^{\prime}$ are $T$-siblings, it follows that the rooted subtrees $T(v)$ and $T(v^{\prime})$ have exactly the same neighbors in the set $V(Z)\setminus(V(T(v))\cup V(T(v^{\prime})))$, that is,
@@
Equation 8 implies that, up to i1i-1, vertex $v^{\prime}$ has been following exactly the same moves in $\sigma^{\prime}$ as the ones followed by vertex $v$ in $\sigma$. Thus, the fact that $vw\in E(T_{i})$ implies that $v^{\prime}w\in E(T^{\prime}_{i})$, and the claim follows.
@@
Since ${\sf root}(Z)\in M$ and $v\notin M$, the definition of the marking algorithm implies that there exist a vertex $v^{\star}\in{\sf ancestors}(Z,v)$ such that all vertices in $T(v^{\star})$ are non-marked, and a marked $T$-sibling $v^{\prime}$ of $v^{\star}$ with $\tau(T,Z,v^{\star})=\tau(T,Z,v^{\prime})$. Moreover, the fact that $M$ is defined by recursively marking up to $k+1$ vertices of each type implies, together with 4.14 and the fact that the desired vertex $v^{\star}$ is non-marked, imply that we can choose $v^{\prime}$ such that no vertex in $T(v^{\prime})$ is used by $\sigma$. Finally, we can choose $v^{\star}$ in a unique way as being the vertex closest to $v$ satisfying the above properties.
Figure 9: Illustration of Case 2 in the proof of 4.16. All vertices in $T(v^{\star})$ are non-marked, and (at least) vertex $v$ is used by $\sigma$. No vertex in $T(v^{\prime})$ is used by $\sigma$, and (at least) $v^{\prime}$ and all vertices in the image of $\rho$, including $\rho(v)$, are marked (in red). The sets $U_{i,\tau}$ of vertices used by $\sigma$ in $T(v^{\star})\cap Z$ are depicted with squares, as well as their images in $T(v^{\prime})\cap Z$ via the representative function $\rho$.
@@
Assuming that the sets $U^{\prime}_{i,\tau}$ exist, it is clear that $\rho$ is injective. Hence, we shall prove that for every $i=0,\ldots,{\sf dist}_{T}(v^{\star},v)$ and every type $\tau$ occurring in a vertex in $U_{i}$, there exists a set $U^{\prime}_{i,\tau}\subseteq V(T(v^{\prime}))\cap Z$ of marked vertices of type $\tau$ with $|U^{\prime}_{i,\tau}|=|U_{i,\tau}|$. We proceed by induction on $i$. For $i=0$, if $U_{0}\neq\emptyset$, then $U_{0}=\{v^{\star}\}$ and the only type occurring in $U_{0}$ is $\tau(T,Z,v^{\star})=:\tau$. Thus, we can just take $U^{\prime}_{0,\tau}=\{v^{\prime}\}$, where $v^{\prime}$ is the vertex given by 6, so that $\rho(v^{\star})=v^{\prime}$. Note that it is not necessarily the case that vertex $v^{\star}$ is used by $\sigma$, which is equivalent, by the above definitions, to $U_{0}\neq\emptyset$.
We now prove the statement for any $i\geq 1$, using the recursive definition of type (cf. 4.9). We may assume that $U_{i}\neq\emptyset$, as otherwise there is nothing to prove. The facts that $\tau(T,Z,v^{\star})=\tau(T,Z,v^{\prime})$ and that $v^{\prime}$ is marked imply that, for every $i\in[{\sf dist}_{T}(v^{\star},v)]$ and every type $\tau$, the following holds:
@@
  • Suppose first that none of $u,w$ belongs to $T(v^{\star})$. It is not difficult to verify that if $uw\in E(T_{i-1})$, then $uw\in E(T^{\prime}_{i-1})$ as well, where $T^{\prime}_{i-1}$ is the tree obtained from $T$ by applying the first i1i-1 rotations of $\sigma^{\prime}$. Indeed, the fact that $\tau(T,Z,v^{\star})=\tau(T,Z,v^{\prime})$ implies that the existence of such an edge with both endpoints outside $T(v^{\star})$ is preserved when replacing the vertices in $U$ with their representatives.
@@
    Assume now inductively that the edges to be rotated exist up to i1i-1, and suppose that $e_{i+1}=uw$ for some $u\in V(T(v^{\star}))$ and $w\notin V(T(v^{\star}))$. The fact that $\tau(T,Z,u)=\tau(T,Z,\rho(u))$ implies that ${\sf trace}(T,Z,u)={\sf trace}(T,Z,\rho(u))$, which implies in particular that $T(u)$ and $T(\rho(u))$ have exactly the same neighbors in the set $V(Z)\setminus(V(T(v^{\star}))\cup V(T(v^{\prime})))$, that is,
@@
    The fact that $\tau(T,Z,u)=\tau(T,Z,\rho(u))$ and Equation 9 imply that, up to i1i-1, vertex $\rho(u)$ has been following the same moves within $T(v^{\prime})$ and $V(Z)\setminus(V(T(v^{\star}))\cup V(T(v^{\prime})))$ in $\sigma^{\prime}$ as the ones followed by vertex $u$ within $T(v^{\star})$ and $V(Z)\setminus(V(T(v^{\star}))\cup V(T(v^{\prime})))$ in $\sigma$. Thus, the fact that $uw\in E(T_{i})$ implies that $\rho(u)w\in E(T^{\prime}_{i})$, and the claim follows.
@@
By 9, $\sigma^{\prime}$ is an $\ell$-rotation sequence from $T$ to $T^{\prime}$, and it uses strictly more marked vertices than $\sigma$, because no vertex of $T(v^{\star})$ is marked (by the conditions in 6), and within $T(v^{\prime})$ there is at least one marked vertex used by $\sigma^{\prime}$, namely $\rho(v)$ (cf. Figure 9). This concludes Case 2. O